OpenStudy (softballgirl372015):

Help!! One angle of a triangle measures 65º. The side opposite this angle has length 10. If the area of the triangle is 20, find the perimeter.

OpenStudy (espex):

This one will likely require collecting all the equations you know about a triangle and its area/angles/perimeter and solve the long way.

OpenStudy (anonymous):

Do you know what type of triangle, such as right or equilateral? That would help a lot.

OpenStudy (softballgirl372015):

it was not given the type of triangle.

OpenStudy (anonymous):

is there a picture to go with it?

OpenStudy (softballgirl372015):

no :(

OpenStudy (anonymous):

Hmm. A = 1/2bh 20 = 1/2bh 20 = 1/2 (10)h

OpenStudy (softballgirl372015):

The answer is perimeter=25, but I do not know how to get to this answer.

OpenStudy (anonymous):

Solve for h then you will have 2 numbers toward the perimeter

OpenStudy (anonymous):

That's really helpful! So whatever number you get for h. Do 25 - 10 - h

OpenStudy (softballgirl372015):

I tried doing it that way, but assumes that it is either a right or isosceles triangle. And I did not arrive at the answer, perimeter=25.

OpenStudy (softballgirl372015):

For this problem I have to show my work towards the answer of perimeter=25. That information is not given in the problem.

OpenStudy (anonymous):

25 - 10 - 4 = 11 is what I got

OpenStudy (softballgirl372015):

Okay, thank you for trying to help me. I will have to ask my teacher tomorrow.